3x-4y=12 intersect an ellipse x^2/25+y^2/16=1  at P and Q find the point of intersection of tagents at P amd Q.Please. send me a detailed solution.Thankyou
 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
This line will act as a chord of contact for point (h, k)
 
Now
Chord of contact is T = 0
 
 
hx /25  + ky/16 = 1
 
Now compare this to
 
3x - 4y = 12
 
h/25 = 3/12
h = 75/12
And
 
k/16 = -4/12
k = -16/3
 
Hope it helps